OpenStudy (anonymous):

what percent of 850 is 238. use a porpotion to answer the question

OpenStudy (unklerhaukus):

238:850 =28:100

OpenStudy (anonymous):

can you please explain :)

hero (hero):

is/of = %/100 238/850 = x/100

hero (hero):

Solve for x

hero (hero):

23800 = 850x\ 23800/850 = x

hero (hero):

x = 28

hero (hero):

So 28%

OpenStudy (anonymous):

UncleRhaukus got it right. He just scaled to a scale of 100 This simply means that if 850 is 100% of something then 238 would be 28% of that same thing

OpenStudy (anonymous):

thanks everybody

OpenStudy (anonymous):

Hero has just done the math :) so teh problem looks like it's been solved.

hero (hero):

Yeah, showing the steps is very helpful

OpenStudy (anonymous):

thanks again